OpenStudy (anonymous):

the following angles are complementary angles. determine the measures of 5x+ 6 and x

OpenStudy (asnaseer):

do you know what "complementary angles" means?

OpenStudy (anonymous):

Yes 90 degrees

OpenStudy (asnaseer):

correct. so you are given two angles: 5x+6 and x

OpenStudy (asnaseer):

and you are told that they are complementary

OpenStudy (asnaseer):

so what can you say about them?

OpenStudy (anonymous):

So I have to get them to add up to 90 right, but how do I figure out what numbers to plug into the solution?

OpenStudy (asnaseer):

you are correct - so just form an equation - can you do that?

OpenStudy (ghazi):

complementary angles mean..sum of theirs = 90 degrees

OpenStudy (ghazi):

5x+6+x=90

OpenStudy (ghazi):

solve for x

OpenStudy (anonymous):

So how do I determine what number X is?

OpenStudy (asnaseer):

first try to simplify the sum

OpenStudy (asnaseer):

it looks like ghazi has done the initial work for you

OpenStudy (asnaseer):

so you have: 5x+6+x=90 add up the "like" terms here in order to simplify this - can you do that?

OpenStudy (anonymous):

5x^2+6=90?

OpenStudy (asnaseer):

\[5x + x \ne5x^2\]\[5x\times x=5x^2\]

OpenStudy (asnaseer):

you are adding here, not multiplying

OpenStudy (asnaseer):

think of this as "five apples plus one apple"

OpenStudy (anonymous):

I'm confused

OpenStudy (asnaseer):

what is: 5x + x = ?

OpenStudy (asnaseer):

in words this might be phrased as: what is 5 lots of x plus another x equal to

OpenStudy (anonymous):

5x^2

OpenStudy (asnaseer):

no - that is the result of multiplying 5x by x

OpenStudy (asnaseer):

what is 5 lots of x plus another x equal to?

OpenStudy (anonymous):

6x

OpenStudy (asnaseer):

correct, so now, going back to the original problem, we had: 5x + 6 + x = 90 which you can simplify down to: 6x + 6 = 90

OpenStudy (asnaseer):

make sense so far?

OpenStudy (anonymous):

Okay

OpenStudy (asnaseer):

the next step would be to subtract 6 from both sides of the equation. what do you get when you do that?

OpenStudy (anonymous):

6x=84?

OpenStudy (asnaseer):

perfect! now you just need to divide both sides by 6

OpenStudy (anonymous):

14

OpenStudy (anonymous):

So x=14?

OpenStudy (asnaseer):

thats it! well done! you can always double check to see if this is correct: 5x + 6 = 5*14 + 6 = 70 + 6 = 76 x = 14 so: 5x+6 + x = 76 + 14 = 90

OpenStudy (anonymous):

So one angle is 76 and the other is 14 correct?

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

Thank you!

OpenStudy (asnaseer):

yw :)